OpenStudy (anonymous):

Show all work to simplify [(1/x)+(3/8)=(1/2)]

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large{1 \over x} + {3 \over 8} = {1 \over 2} }\) \( \color{Black}{\Rightarrow \Large{1 \over x} = {1 \over 2} - {3 \over 8} }\) Do you understand it till this step?

OpenStudy (anonymous):

yeah you just moved it around. i get that

Parth (parthkohli):

Very good. Now, what is \(\Large {1 \over 2} - {3 \over 8}?\)

OpenStudy (anonymous):

1/8

Parth (parthkohli):

Correct!

Parth (parthkohli):

Now \(\Large {1 \over x} \normalsize = \Large{1 \over 8}\). Can you tell me what \(x\) is?

OpenStudy (anonymous):

i think so, hold on

OpenStudy (anonymous):

actually can you explain to me how to solve that. im confused:/

Parth (parthkohli):

You just multiply both sides by \(x\), and then multiply both sides by \(8\).

OpenStudy (anonymous):

(1/x)+(3/8)=(1/2) (1/x)+(3/8)-(3/8)=(1/2)-(3/8) (1/x)=(4/8)-(3/8) (1/x)=(1/8) x=8

Parth (parthkohli):

So: \( \color{Black}{\Rightarrow 1 = \Large{x \over 8} }\)

Parth (parthkohli):

Now if you actually try to multiply 8 to both sides, you'd get \(x = 8 \times 1\)

OpenStudy (anonymous):

ahh i think i understand now. i neeed help with another problem tho..

Parth (parthkohli):

\(8 \times 1 = 8\), so \(x = 8\).

Parth (parthkohli):

@joemost Sure, post it in a new one :)

OpenStudy (anonymous):

just posted it